donna invested her $33000.00 bonus and received a total of $970.00 in interest after one year. if part of the money returned 4% and the remainder 2.25%, then how much did she invest at each rate?
Found 2 solutions by Mathtut, edjones:
Answer by Mathtut(3670)

well there are two amounts we are dealing with here that total $33000. so we will call one amount x and the other 33000-x. the total of the two principal sums times there various interest rates will equal the interest earned.
so we have .04x+.0225(33000-x)=970 distribute the 2nd term and we get
.04x+742.50-.0225x=970 simply to .0125x=227.5 x=$18200 ... $33000-x=$14800
so Donna invested $18,200 at 4% and $14,200 at 2.25%

Answer by edjones(7311)

Let x be the amt invested at 4%.
.04x+.0225(33000-x)=970
.04x+742.50-.0225x=970
.0175x+742.50=970
.0175x=227.5
x=$13,000.
33000-13000=$20,000
.
Ed